The fitness of 514 adults (aged 25 to 35) were sampled and their heart rates measured immediately after exercise. This sample of adults was found to have a mean of 1427 beats (per 10 minutes) with a standard deviation of 325 beats (per 10 minutes).

Construct a two-tailed 95% confidence interval for the true mean heart beat of all such adults.

Determine the sample size that would be necessary to obtain an interval with width (at most) 50 for a confidence level of 95%.

Let assume the sample is from a normal population, and sample size is 30 instead of 514. Construct a

99% upper tail confidence interval for the true mean heart beat of all such adults.
Answer #1

Mean= 1417

Std= 325

n= 514

a) 95% Confidence interval: t_{alpha /2,n-1}=t_{0.05,513}=1.964599

CI: ar{X}pm t*(S/sqrt{n})=>1427pm 1.964599*(325/sqrt{514})

CI (1398.837,1455.163)

b) ME= 50

ME=2* Z*(S/sqrt{n})=> n= (rac{2*Z*S}{ME})^2= (rac{2*1.96*325}{50})^2= 649

c) n=30

99% Confidence interval: t_{alpha ,n-1}=t_{0.01,29}=2.462021

UCI:ar{X}+ t*(S/sqrt{n})=>1427+ 2.462021(325/sqrt{30})

UCI:1573.088